# Artificial intelligence for science: The easy and hard problems

### v1

@@ Abstract
Several major conceptual breakthroughs led to the “protein folding problem.” The discovery that proteins are linear chains of amino acids goes back to the seminal sequencing of insulin by Frederick Sanger in 1951 [32], based on the isolation and recursive extraction of hydrolyzed protein fragments using various media and electrical currents. Evidence that the overall 3D structure of proteins was important for their function, rather than the identity of individual amino acids, [^4] The notable exception to this proposition is the identity of certain amino acids in the active site of enzymes. came from X-ray crystallography of oxygen-carrying proteins [33, 34], the structural effects of natural and artificial variation of amino acids [35], and catalytic-rate analyses with different cellular conditions, substrates, and inhibitors [36, 37].
The third development was the specification of the protein folding problem, primarily by Christian Anfinsen Jr. [38], who found that ribonuclease A would lose its enzymatic activity in artificial conditions and recover it when physiological conditions were re-established. This led to the “thermodynamic hypothesis” that the correctly folded protein occupied the minimum free-energy state in its natural cellular environment, and provided evidence against the competing hypothesis that proteins folded sequentially as they were synthesized. [^5] There were data that this theory did not apply to—for example, proteins that required enzymatic modification to renature; or assistance during folding by “chaperone” proteins. The remaining step was to characterize the physical process by which the protein folded.
@@
One of the most successful recent discovery algorithms is AlphaFold2 [11], which predicts the 3D structure of a protein given its 1D amino-acid sequence (see Box 3). When it was released, AlphaFold2 brought the average molecular deviation for a protein down from 0.3 to 0.1 nanometers, which was precise enough for biologists to make use of.
AlphaFold2’s input is a multiple sequence alignment (MSA), which augments the protein of interest’s 1D amino-acid sequence with additional rows containing similar amino-acid sequences from existing databases. If any of the MSA sequences have already had structures derived, 2D distograms of the pairwise distance between residues and a sequence of torsion angles between adjacent amino acid residues are added to the inputs.
AlphaFold2 outputs a set of atomic co-ordinates, a confidence score in each residue’s position, torsion angles between adjacent amino-acid backbones, the 2D distogram between residues, and a prediction of any masked parts of the MSA. The objective function during training contains a loss term for each of these representations, with the most important components penalizing the 3D deviations of heavy atoms in the amino-acid chain. The loss function during “fine-tuning” contains all of these terms, plus two extra terms that penalize the final structure for violating physical constraints.
@@
AlphaFold2 uses complex heuristics to solve this optimization problem, based on a great deal of biological and engineering knowledge. At a high level, the Evoformer module learns increasingly rich and abstract representations of the 1D primary structure and 2D distogram that the Structure module uses to build a 3D model of the protein. The network is trained end-to-end, meaning all operations are differentiable and the loss signal from the final 3D positions is back-propagated to inform the update of neural networks weights in all operations after the input.
The main biological insight behind the Evoformer module is that information about the 3D protein structure can be derived by comparing its primary sequence with the sequences of similar proteins in different organisms. Some of these sequences might have had their structures discovered experimentally, which can be used directly in the 2D distance representation. But even when no related structure exists, significant covariation of residues in two different positions across multiple organisms is an indication that they are close in 3D space. These 3D dependencies might be quite far away in the 1D representation (the primary sequence), so the inductive bias of attention, which can model longer dependencies [39], is more suitable than other deep learning methods. When a particular structure is available, it biases the attention mechanism to learn similar representations for amino acids that are physically close together, and when it is not, the information flows the other way, with the covariance between amino acid positions used to infer the 2D distances.
The Structure module uses the Evoformer’s final representations to iteratively move rigid frames representing each amino-acid residue as close as possible to their ground truth cognates. After residues have been aligned in the main training cycle, refinement steps add the amino acids’ side chains and alter their positions during fine-tuning such that inter-atomic bonds take physically plausible values and no side-chains overlap.
@@
AlphaFold2: Analysis
AlphaFold2’s success comes in large part from the engineering choice of problem statement. In particular, it does not solve the original “problem” of protein folding, the time-evolving movement of the polypeptide chain from 1D denatured to 3D functional state. The authors relax the physical requirements to define a new, related problem: The prediction of a final folded state given the 1D sequence.
This choice was motivated by an abundance of sequence data, which can be used for the new, but not the old, problem. [^6] The MSA can be generated using off-the-shelf tools from genomics, which now have very large sequence datasets. It was also based on a suite of models with the ability to use that information to solve the relaxed problem formulation—deep neural networks with attention mechanisms. The positive consequence of this choice is that some requirements of a solution to the original problem are met—we can predict the structure of hydrophilic proteins with lots of analogous evolutionary sequences well. The negative consequence is that we don’t have a model of the folding dynamics which can make good predictions of the structures of orphan molecules like antibodies, lipophilic molecules with no experimentally derived homologous structures, or the effect of a new mutation or ion on the final folded structure.
The choice of inputs and outputs reflect the new problem. Evolutionary correlations have been used for some time to make arguments about folded structures and function [40], but do not obviously inform folding dynamics. And, the pair-wise distogram input has been chosen as a primitive because this is the form required by the algorithm that will be used upstream (attention). AlphaFold2’s intermediate and output representations and transformations constitute a highly distributed theory. The most important engineering choice, in light of defining the new problem, is to partition training into a free optimization period over the residue gas representation, plays to the strengths of deep learning in iterative representation learning and stochastic gradient descent, followed by a fine-tuning stage where physical constraints are met.
Although the scale and complexity of natural biological systems warrants different types of theory and strategies of investigation, including decomposition and localization [41], there are also important commonalities—including the use of ontologies [42] and imagistic intermediate models [37, 40]. Conversely, Lavoisier and Maxwell were often tempted to re-specify constraints based on abundant data and a powerful method nearby. [^7] Lavoisier’s use of the Hales’ apparatus and burning lens is a good example of this. We would like AI scientists that can likewise recognize when progress has been slow on a particular problem, but adjacent sources of data and powerful models promise fulfilment on an intersecting set of desiderata. We would also like them to recognize when and how to gather more useful data when there is a mismatch with the use case—as recent improvements on using AlphaFold2 to predict human structures have done.
@@
An important and elusive feature of problem specification is that it is not a data modeling problem. The selection of what to model and and what constraints to condition on are antecedent to any data modeling problem. It is also not reducible to a representation learning problem, in the sense of figuring out how raw sensory input maps to abstract representations. Of course, that problem also needs to be solved, but first the scientist needs to know what problems the representations are being used to solve.
@@
A research program for attacking the hard problem should begin with the cognitive science of science [43], focusing on the understudied subjective, creative aspects discussed above and how they interact with the objective aspects of problem solving. However, this presents two immediate challenges. First, how can we gain the conceptual background necessary to understand scientists’ innovations in a short enough time to iterate meaningful research? For most graduate students, arriving at the point where they can begin to generate meaningful and achievable problems within their field takes 2-5 years of dedicated higher education. Second, how can we gather enough results to make statistically robust arguments for any individual problem? There was, of course, only one Antoine Lavoisier.
Cognitive-historical analyses are one approach to deriving such insights [27, 42, 44]. In this methodology, modern cognitive theories are used to build hypotheses about how the scientists were thinking. Historical data can add content or temper these theories, and historical analyses and techniques can be used to make the retrospective analyses relatively unbiased, robust to historical contingencies, and generalizable to new contingencies. At the birth of a modern scientific field, the concepts and measurements are relatively undifferentiated, and can be acquired quickly. They are also, necessarily, edge-cases of creativity, where one or a group of scientists broke away from the normal tradition.
Spending time observing scientists’ behaviors in modern operating laboratories is another way to increase the amount of data available for cognitive scientists to build theories about problem specification. For instance, the construction of intermediate in-vitro models as sources of analogy has been hypothesized to explain the success of scientific research practices in biochemistry [45]. Studies of scientific collaborations between people from different fields or methodological backgrounds have emphasized the importance of visual aids and hand gestures in providing explanations [46].
The birth of the internet and online crowdsourcing platforms has allowed cognitive psychologists to scale up their studies to online experiments where the behaviors of very large numbers of computer-literate participants can be tested [47]. We can use our insights from cognitive-historical analyses and laboratory observations to design prospective tests of the key computational principles underlying the construction of problems and other aspects of discovery. Indeed, related studies already provide strong evidence that humans construct simplified mental representations to plan [48], but have not been extended to the less well-defined problem settings in science. Another rich set of problems come from tests of physical reasoning, in which previous laboratory-based work has identified iterative model-based revision of problem statements as a critical part of deriving a successful scientific solution [49, 50].
@@
Once we understand what human scientists are doing with enough precision that we can formalize their activities, we can try to leverage these insights to build scalable AI scientists. At least initially, it is unlikely that these will be standalone systems, but rather more like research assistants or first-year grad students: Curious agents with some technical competence but in need of expert guidance. This guidance can come in the form of natural language instruction, reading curricula, and demonstrations. The growth of models beyond this requires the examination and emulation of the communal aspects of science and related cultural institutions. Lab meetings, conferences, and presentations and discussions are ultimately the place where judgements on the quality of a scientific problem are made.
@@
The use of natural language processing for scientific discovery is at the heart of the recently proposed “AI Scientist” [51], which autonomously updates machine-learning (ML) code in order to generate scientific papers. In its inner loop the AI Scientist is given access to the training, testing, and visualization code for a simple ML model and dataset, along with several suggestions of innovative changes to the code and the overall objective of reducing the model’s loss on held-out data. Its outer loop requires that it generate a range of ideas in natural language format, check their novelty using the internet, apply several ideas, write a ML paper for each of the ideas that ran successfully, review the paper, then update the paper. The proposed system comprises a carefully designed interface of language models, prompting schemes, a coding assistant, and templates for papers and conference guidelines. From the examples presented in [51], the innovative ideas that the AI Scientist generates are mostly decisions to split variables or processing pathways, add new model components or training metrics based on previously successful strategies in the literature, and combine any of the above that improve the final loss. A particularly impressive part of the work is the ability to implement these high-level conceptual changes in the code example, including producing useful visualizations.
@@
Whether this system and its successors can produce radically innovative discoveries remains to be seen. Do such systems replicate human strategies such as ontologically guided constraint respecification, producing and modifying intermediate models, and re-specifying the problem based on knowledge of adjacent rich sources of data and available models? Natural language is certainly capable of capturing some aspects of the ontological structure of knowledge, and multimodal models should be able to create and maintain imagistic intermediate models of the scientific phenomenon.
@@
On the other hand, many scientific developments, including those we have characterized above, come from a reflective consideration of either how to alter model constraints to capture anomalous data [52, 42], or where an alteration of model constraints affects the domain, borne out over a course of successive investigations [22, 27]. Whether current models that decouple in-context and weight-based learning can capture this type of reflective continual learning and selective conceptual respecification will require further investigation [53, 54]. For now, humans remain the only intelligent system capable of solving the hard problem. We still have much to learn about building AI scientists by studying ourselves.
@@ Acknowledgments
We are grateful to Nancy Nersessian for helpful discussions. This work was supported by the Kempner Institute for the Study of Natural and Artificial Intelligence, and by the Schmidt Science Polymath Program.

### v2

@@ Abstract
Like Lavoisier, Maxwell was guided in this process of abstraction by ontological knowledge about the structure of different physical and mathematical systems, which also helped him sequentially assemble and modify the mathematical expressions underlying the model. Perhaps these idealized models played a role in Lavoisier’s early investigations, albeit in a simpler form involving crude movements of air and changes of weight. What is clear is that this process is not captured by systems like AI Feynman, which are given the problem variables from the mature idealized model, and lack the flexibility to alter their own conceptual systems—analogous to providing the symbol $i$ before deriving a general solution to the problem of polynomial root finding.
The kinds of laws AI Feynman can derive are also limited by its processing steps. This is motivated by an analysis of common characteristics of physical laws—they contain variables with units, low-degree polynomial structure, compositionality, smoothness, symmetry, and separability. But again, these constraints arose out of analysis of the existing laws of physics, and provide constraints that restrict the subsequent class of models in an inflexible manner—Maxwell invented dimensional analysis to help solve difficult physics problems.
And, once again, there there is also the question of which datapoints are chosen. For mystery table 5, the data are not taken from systems far from the scientist or near large masses, where the behavior of light (its speed or deflection, respectively) needs to be taken into account. Recognizing and adjusting for these factors were essential parts of proving the theory and then taking it forward.
@@
Several major conceptual breakthroughs led to the “protein-folding problem.” Frederick Sanger discovered that proteins are linear chains of amino acids based on the isolation and recursive extraction of hydrolyzed fragments of insulin using various media and electrical currents [25]. Evidence that the function of proteins depended on their 3D structure, rather than the identity of individual amino acids, came from X-ray crystallography [26], the structural effects of natural and artificial variation of amino acids [27], and catalytic-rate analyses with different cellular conditions, substrates, and inhibitors [28]. But it was Christian Anfinsen Jr. that put forward what became known as “Anfinsen’s dogma”—that under physiological conditions the primary sequence itself, and no other factors, determined the 3D structure [29].
@@
One of the most successful recent discovery algorithms is AlphaFold2 [7], which is a large neural network model that predicts the 3D structure of a protein given its 1D amino-acid sequence (see Box 3). When it was released, AlphaFold2 brought the average molecular deviation for a protein down from 0.3 to 0.1 nanometers, which was finally precise enough for biologists to make use of. AlphaFold2 uses complex heuristics to solve this optimization problem, based on a great deal of biological and engineering knowledge. At a high level, its Evoformer module learns increasingly rich and abstract representations of the 1D primary structure and and a 2D matrix of inter-atomic distances between residues, and its Structure module uses these representations to build a 3D model of the protein. The network is trained end-to-end, meaning all operations are differentiable and the loss signal from the final 3D positions is back-propagated to inform the update of neural networks weights in all operations after the input.
AlphaFold2’s input is a multiple sequence alignment (MSA), which augments the protein of interest’s 1D amino-acid sequence with additional rows containing similar amino-acid sequences from existing databases. If any of the MSA’s sequences have already had structures derived, 2D distograms of the pairwise distance between residues and a sequence of torsion angles between adjacent amino acid residues are added to the inputs. AlphaFold2 outputs a set of atomic co-ordinates, a confidence score in each residue’s position, torsion angles between adjacent amino-acid backbones, the 2D distogram between residues, and a prediction of any masked parts of the MSA. The objective function during training contains a loss term for each of these representations, with the most important components penalizing the 3D deviations of heavy atoms in the amino-acid chain. The loss function during “fine-tuning” contains all of these terms, plus two extra terms that penalize the final structure for violating physical constraints.
Anfinsen
AlphaFold2’s success comes in large part from the engineering choice of problem statement. In particular, it does not solve the original “problem” of protein folding. Anfinsen and his colleagues began by studying bovine ribonuclease, a protein that could be isolated in abundance using the techniques of the time. Ribonuclease has a functional structure that depends on the (single) correct formation of four disulphide bonds (out of a possible 105 combinations between eight sulfhydryl groups). When these were broken, then allowed to reform in denaturing conditions (containing a high concentration of urea), an inactive mixture of all the possible configurations was generated. In physiological conditions, however, the mixture was converted to the native ribonuclease. Anfinsen and colleagues moved on to study Staphylococcal nuclease, which does not depend on disulphide bonds for its 3D functional structure, analyzing the folding and functionality of various sub-sections of the molecule as well as the temporal- and pH- dynamics of the rather sharp phase transition between inactive and active molecules and the initial few folding events in renaturation. The cumulation of these events was the “thermodynamic hypothesis”—that the correctly folded protein occupied the minimum free-energy state in its natural cellular environment [29].
@@
However, the task left facing scientists working on the protein-folding problem was daunting: Accounting for the full physical process by which a protein of $N$ amino acids assumed one of its $\sim 8^{N}$ possible conformations. Part of the genius of Alphafold2 was that its architects recognized that this problem did not need to be solved in order to make substantial impact in biology, and they relaxed its constraints to define a new, related problem: The prediction of a final folded state given the 1D sequence, leaving behind the requirement to model the time-evolving movement of the polypeptide chain from 1D denatured to 3D functional state. This respecification allowed them to bring in an abundance of sequence data, which can be used for the new, but not the old, problem. Evolutionary correlations have been used for some time to make arguments about folded structures and function [30], but do not obviously inform folding dynamics. It also allowed them to bring in the very flexible learning algorithms from modern deep learning. In particular, attention-based mechanisms model longer-range dependencies over 1D sequences, making use of the 2D distogram representations, and training could be split into a free optimization period, where physical constraints can be violated during iterative representation learning and stochastic gradient descent over the residue gas representation, followed by a bespoke fine-tuning stage where physical constraints were met.
The positive consequence of this choice is that some requirements of a solution to the original problem are met—we can predict the structure of hydrophilic proteins with lots of analogous evolutionary sequences well. The negative consequence is that we don’t have a model of the folding dynamics which can make good predictions of the structures of orphan molecules like antibodies, lipophilic molecules with no experimentally derived homologous structures, or the effect of a new mutation or ion on the final folded structure.
Although the scale and complexity of natural biological systems warrants different types of theory and strategies of investigation, including decomposition and localization [31], there are also important commonalities—including the use of ontologies [32] and imagistic intermediate models [20, 30]. Conversely, Lavoisier and Maxwell often re-specified problem constraints to bring in abundant data and a powerful method nearby.
@@
An important and elusive feature of problem specification is that it is not a data modeling problem. The selection of what to model and and what constraints to condition on are antecedent to any data modeling problem. It is also not reducible to a representation learning problem, in the sense of figuring out how raw sensory input maps to abstract representations. Of course, that problem also needs to be solved, but first the scientist needs to know what problems the representations are being used to solve. Are these conceptual breakthroughs just patterns that can be discovered with a sufficiently powerful pattern recognition system? In a sense yes, but before that can happen, something has to tell the pattern recognition system what kind of patterns are interesting, important, and useful. What problem is the pattern-recognition system designed to solve, and where does this come from?
@@
A research program for attacking the hard problem should begin with the cognitive science of science [33], focusing on the understudied subjective, creative aspects discussed above and how they interact with the objective aspects of problem solving. In the study of science, there is a natural continuum through rare but large breakaway conceptual developments, best studied through cognitive-historical analyses [20, 32, 34]; daily scientific problem solving, studied through observation of practising labs [35]; and large datasets of decisions in simplified problem-solving settings in online naturalistic games [36]. We can tests any hypotheses about science that we derive from these analyses in large-scale online behavioural experiments, where, for example, related studies already provide evidence that humans construct simplified mental representations to plan [37], and that iterative model-based revision of problem statements as a critical part of deriving a successful scientific solution [38].
@@
Once we understand what human scientists are doing with enough precision that we can formalize their activities, we can try to leverage these insights to build scalable AI scientists. At least initially, it is unlikely that these will be stand-alone systems, but rather more like research assistants or first-year grad students: Curious agents with some technical competence but in need of expert guidance. This guidance can come in the form of natural language instruction, reading curricula, and demonstrations. The growth of models beyond this requires the examination and emulation of the communal aspects of science and related cultural institutions. Lab meetings, conferences, and presentations and discussions are ultimately the place where judgements on the quality of a scientific problem are made.
@@
The use of natural language processing for scientific discovery is at the heart of the recently proposed “AI Scientist” [39], arguably the first artificial scientist that addresses the hard problem. The algorithm itself is a carefully designed system of large language models (LLMs), prompting schemes, a coding assistant, and templates for papers and conference guidelines that autonomously updates machine-learning (ML) code in order to generate scientific papers. In its inner loop the AI Scientist is given access to the training, testing, and visualization code for a simple ML model and dataset, along with several suggestions of innovative changes to the code and the overall objective of reducing the model’s loss on held-out data. Its outer loop requires that it generate a range of ideas in natural language format, check their novelty using the internet, apply several ideas, write a ML paper for each of the ideas that ran successfully, review the paper, then update the paper. The proposed system comprises a carefully designed interface of language models, prompting schemes, a coding assistant, and templates for papers and conference guidelines. From the examples presented, the innovative ideas that the AI Scientist generates are mostly decisions to split variables or processing pathways, add new model components or training metrics based on previously successful strategies in the literature, and combine any of the above that improve the final loss. A particularly impressive part of the work is the ability to implement these high-level conceptual changes in the code example, including producing useful visualizations.
@@
Whether this system and its successors can produce radically innovative discoveries remains to be seen. Do such systems replicate human strategies such as ontologically guided constraint respecification, producing and modifying intermediate models, and re-specifying the problem based on knowledge of adjacent rich sources of data and available models? Natural language is certainly capable of capturing some aspects of the ontological structure of knowledge, and multimodal models should be able to create and maintain imagistic intermediate models of the scientific phenomenon. We would like AI scientists that can likewise recognize when progress has been slow on a particular problem, but adjacent sources of data and powerful models promise fulfilment on an intersecting set of desiderata. We would also like them to recognize when and how to gather more useful data when there is a mismatch with the use case—as recent improvements on using AlphaFold2 to predict human structures have done.
@@
On the other hand, many scientific developments, including those we have characterized above, come from a reflective consideration of either how to alter model constraints to capture anomalous data [40, 32], or where an alteration of model constraints affects the domain, borne out over a course of successive investigations [17, 20]. Whether current large language models can capture this type of reflective continual learning and selective conceptual respecification will require further investigation [41, 42]. For now, humans remain the only intelligent system capable of solving the hard problem. We still have much to learn about building AI scientists by studying ourselves.
@@ Acknowledgments
We are grateful to Nancy Nersessian, Melanie Mitchell, Giovanni Pezzulo, Frank Keil, Skyler Wang, George Davis-Smith, John Clement, and Tom Griffiths for helpful discussions. This work was supported by the Kempner Institute for the Study of Natural and Artificial Intelligence, and by the Schmidt Sciences Polymath Program.